Synovec Co. is growing quickly. Dividends are expected to grow at a rate of 2.31 percent for the next three years, with the growth rate falling off to a constant 5.46 percent thereafter. If the required return is 8.77 percent and the company just paid a dividend of $1.22, what is the current share price?
Answer #1

D1=(1.22*1.0231)=1.248182

D2=(1.248182*1.0231)=1.277015004

D3=(1.277015004*1.0231)=1.306514051

Value after year 3=(D3*Growth Rate)/(Required rate-Growth rate)

=(1.306514051*1.0546)/(0.0877-0.0546)

=$41.6268797

Hence current share price=Future dividends and value*Present value of discounting factor(rate%,time period)

=1.248182/1.0877+1.277015004/1.0877^2+1.306514051/1.0877^3+$41.6268797/1.0877^3

=$35.59(Approx).